OpenStudy (anonymous):

Carrie started with 12 ounces of water when she began her walk. When she was finished walking she had 2 1/4 ounces of water left. How much did she drinK?

OpenStudy (anonymous):

Isn't it just 12 - 2 1/4 ?

OpenStudy (anonymous):

\[12-\frac{ 9 }{ 4 }= \frac{ 48 }{ 4 }-\frac{ 9 }{ 4 }= \frac{ 39 }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ 39 }{ 4 }= 9\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

Carrie drank 9 3/4 ounces of water.